OpenStudy (dls):

If 2 numbers A and B are selected from the set of first 100 natural numbers then in how many ways,

OpenStudy (dls):

\(a^2-b^2\) is divisible by 3?

OpenStudy (dls):

Lets try.. Numbers are of form: 3k,3k+1,3k+2 Squared numbers are of form: \[(3k)^2=9k^2\] \[(3k+1)^2=9k^2+6k+1=3k(3k+2)+1\] \[(3k+2)^2=9k^2+12k+4=3k(3k+4)+4\]

OpenStudy (anonymous):

Remember, \(r\in \{0,1,2\}\)

OpenStudy (dls):

yup

OpenStudy (anonymous):

4 = 3+1

OpenStudy (dls):

what do you mean?

OpenStudy (anonymous):

\[ 4\equiv 1 \pmod 3 \]

OpenStudy (dls):

didn't get what that meant since morning :p

OpenStudy (anonymous):

You don't know what modulus is, but you know what divisibility is?!

OpenStudy (dls):

I know modulus but what are you relating here I'm not getting

OpenStudy (anonymous):

I'm saying that the remainder of \((3k+2)^2\) is \(1\) rather than \(4\).

OpenStudy (dls):

how ? :/

OpenStudy (dls):

okay yes

OpenStudy (dls):

got it 4=1+3 :P

OpenStudy (anonymous):

You're killing me!

OpenStudy (dls):

So squared numbers are of form : 3k,3k+1,3k+1 (do i need to write this twice?)

OpenStudy (anonymous):

Writing it twice can help.

OpenStudy (dls):

Let us consider all the cases now :D 1)A of form 3k,B of form 3k 2)Both of form 3k+2 3)A of form 3k+1,B of 3k+2 4)A of form 3k+2,B of form 3k+1

OpenStudy (anonymous):

No, which combinations of these remainders has remainder 0?

OpenStudy (anonymous):

Now^

OpenStudy (dls):

Did I leave any case?

OpenStudy (anonymous):

There were \(3\times 3\) cases.

OpenStudy (dls):

i am still getting confused :/

OpenStudy (anonymous):

There are \(9\) cases.

OpenStudy (dls):

The question is "a^2-b^2" divisible by 3 where did we use that info?I mean same solution for a^2+b^2 ? why did we compute the squared form of numbers? we dont have 3k+2 in that but then how can we assume it in the cases

OpenStudy (anonymous):

\((3k+2)^2=3k'+1\)

OpenStudy (anonymous):

Okay, so when considering the TOTAL cases, what we consider is how many remainders \(a\) could have (3) and how many \(b\) can have(3). Combine this and you get \(3\times3=9\) cases.

OpenStudy (anonymous):

The next step is to find out which of those \(9\) cases end up being divisible by three.

OpenStudy (dls):

yes

OpenStudy (anonymous):

As it turns out, there are only two possible remainders you can get after squaring: 0 and 1.

OpenStudy (dls):

yes!

OpenStudy (anonymous):

Now, we need to consider the cases: 0+0 = 0 0+1 = 1 1+0 = 1 1+1 = 2

OpenStudy (anonymous):

Now, 1+1=2 isn't divisible by 3 because it doesn't have remainder of 0. But hypothetically if it WAS divisible by 3.. Then it since there are 2 ways to have remainder 1 ( 3k+1, 3k+2 ) Then the ways to get 1+1 is \(2\times2=4\)

OpenStudy (anonymous):

Fortunately the only case divisible by \(3\) is: \(0+0=1\) There is only \(1\times 1=1\) ways for this to happen.

OpenStudy (dls):

thought: will we get same answer if question says a^2 "+" b^2 ?

OpenStudy (anonymous):

Oh wait.. it should be -?

OpenStudy (anonymous):

Yeah, that is a big difference!

OpenStudy (anonymous):

0-0 = 0 1-0 = 1 0-1 = -1 1-1 = 0

OpenStudy (anonymous):

So 0-0 and 1-1 are divisible by 3 in this case.

OpenStudy (anonymous):

0-0 happens \(1\times1=1\) ways 1-1 happens \(2\times2=4\) ways

OpenStudy (anonymous):

So \(5\) out of \(9\) ways!

OpenStudy (dls):

that is what I was thinking :| but \[(3k+1)^2-(3k+2)^2=1+6k-(4+6k)=-3\] if we take A of form 3k+1 and B of form 3k+2

OpenStudy (dls):

-3 :|

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @DLS that is what I was thinking :| but \[(3k+1)^2-(3k+2)^2=1+6k-(4+6k)=-3\] if we take A of form 3k+1 and B of form 3k+2 \(\color{blue}{\text{End of Quote}}\) \[ (3k+1)^2-(3k+2)^2=3k'+1 - (3k'+1)=0 \]

OpenStudy (dls):

what just happened?

OpenStudy (anonymous):

Ugh...

OpenStudy (anonymous):

Okay, so get this: we ONLY need to worry about the REMAINDER

OpenStudy (dls):

yes true so :o

OpenStudy (anonymous):

So why are you worrying about other pellet?

OpenStudy (anonymous):

About other \(\text{pellet}\)?

OpenStudy (anonymous):

About other \(\text{sh} \text{it}\)

OpenStudy (dls):

but it is giving negative remainder .-.

OpenStudy (anonymous):

If it is giving a negative remainder then: 3-r = ?

OpenStudy (anonymous):

3-1 = 2, the remainder is 2

OpenStudy (anonymous):

Or rather remainder -1 becomes remainder 2

OpenStudy (dls):

okay i got evt

OpenStudy (anonymous):

What I would do is this: only consider the cases where a>b

OpenStudy (anonymous):

then double that result for b>a

OpenStudy (dls):

seems a good idea